Here's the question
It costs $350 for a case of 150 jars of modeling clay. It costs $672 for a case of 450 jars of modeling clay.

What is the unit price for a jar in the case of 150 jars?

What is the unit price for a jar in the case of 450 jars?
Which case of jars of modeling clay is the better buy?
Please help me set up the proportion and answer.

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

350/150
672/450
work those out and decide which is smaller

Answer by stanbon(75887)

Here's the question
It costs $350 for a case of 150 jars of modeling clay. It costs $672 for a case of 450 jars of modeling clay.
What is the unit price for a jar in the case of 150 jars?
What is the unit price for a jar in the case of 450 jars?
Which case of jars of modeling clay is the better buy?
Please help me set up the proportion and answer.
---------------
You have two points: (150,350) and (450,672)
slope: (672-350)/(450-150) = 322/300 = 1.073
---------------------------------------------------------
What is the unit price for a jar in the case of 150 jars?
350/150 = $2.333333
-------------------------------------------------------------
What is the unit price for a jar in the case of 450 jars?
672/450 = $1.49333..
-----
Which case of jars of modeling clay is the better buy?
The case with 450 jars is the better buy.
============================
Cheers,
Stan H.